OpenStudy (anonymous):

I get the inputs are at the top but I don't understand how to evaluate functions? Please help!!

OpenStudy (freckles):

Example: If g(x)=x^4-4x+1 then g(5) can be found by replacing x with 5. So g(5)=5^4-4(5)+1 and of course you can simplify this. So g(5)=625-20+1=605+1=606

OpenStudy (freckles):

So for my function g, when my input is 5 my output is 606